OpenStudy (anonymous):

Solve and Check. √x-3=3 a) x = 7 b) x = 5 c) x = 1 d) x = -1

myininaya (myininaya):

\[\sqrt{x-3}=3? or \sqrt{x}-3=3?\]

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

the square root is over the x-3

myininaya (myininaya):

x-3=9 x=9+3=12 dip is right

OpenStudy (anonymous):

then x=12 ,no match..